OpenStudy (anonymous):

Cathy's Caterers is creating a wedding cake by placing one cylindrical cake with a radius of 3 inches on top of another cylindrical cake with a radius of 5 inches. If both of these cakes measure 4 inches in height, what is the volume, in cubic inches, of the entire wedding cake? 1. 136 π 2. 100 π 3. 36 π 4. 34 π 5. 12 π

OpenStudy (anonymous):

volume of cylinder is given by πr^2h where r is radius and h is height

OpenStudy (anonymous):

thus for cake with radius 3 inch volume is π (3^2)*4=36π thus for cake with radius 5 inch volume is π (5^2)*4=100π

OpenStudy (anonymous):

total 100π +36π=136π